OpenStudy (anonymous):

CAL 3 Problem: f(x,y) = sin(xy^2) P=(휋/4,2) southeast Find the directional derivative of f at P in the given direction?

OpenStudy (anonymous):

\[P = (\pi/4,2)\]

OpenStudy (anonymous):

\[f(x,y)=\sin(xy ^{2})\]

OpenStudy (experimentx):

apply gradient ... and just put those values.

OpenStudy (anonymous):

\[gradient?...(\frac{ \partial f }{ \partial x }, \frac{ \partial f }{ \partial y })\]

OpenStudy (experimentx):

yes

OpenStudy (experimentx):

\[ \vec \nabla f = \hat i \frac{\partial f}{\partial x} + \hat j \frac{\partial f}{\partial y}\]

OpenStudy (anonymous):

P\[P = (i,j) ?\]

OpenStudy (experimentx):

hold on ...

OpenStudy (anonymous):

I was referring to the P in the question

OpenStudy (experimentx):



OpenStudy (anonymous):

how'd you get that?

OpenStudy (experimentx):

no ... i didn't read your question properly, later i realized it was directional derivative and the direction was specified.

OpenStudy (anonymous):

ohh SE

OpenStudy (experimentx):